OpenStudy (anonymous):

Find the area of the region enclosed by the graphs y=e^x , y=e^-4x , x=ln(3)

OpenStudy (shubhamsrg):

Can you plot the regions ? (i.e. draw them)

OpenStudy (foolaroundmath):

\(\displaystyle \int_0^3 e^{x}\, dx\) gives the area between the curve \( y = e^x \) and the x-axis. ( 1 in the drawing) \(\displaystyle \int_0^3 e^{-4x} \, dx\) gives the area between the curve \(y = e^{-4x} \) and the x-axis. ( 2 in the drawing)  Subtract the two areas (1 - 2) to get the value of the area enclosed i.e. \(\displaystyle \int_0^3 (e^{x}-e^{-4x})\, dx\)